The angle of a prism is 60° in the angle of minimum deviation is 56.5°, what will be the refractive index of the prism?

a. 1.5
b. 1.7
c. 1.6
d. 1.55
Answer #1

Relation between refractive index of prism and angle of miniumu deviation is given by:

n = [sin ((A + D)/2)]/sin (A/2)

A = Angle of prism = 60 deg

D = Angle of deviation = 56.5 deg

So,

A/2 = 30 deg

(A + D)/2 = (60 + 56.5)/2 = 58.25 deg

So,

n = sin 58.25 deg/sin 30 deg

n = 1.7

Correct option is B.

Let me know if you've any query.